# The Influence of Exclusion Zones on the Coexistence of Predator and Prey with an Allee Effect

### v1

@@ 6.1. Computations of equilibria 1D
    This indicates that population density blows up as $a\to 0^{+}$ while total population approaches a finite limit. This supports the thin-limit analysis of Section 5. In fact, the formulas (54) and (62) allow us to directly calculate the limit of $\overline{V}(a)$ and its derivative as $a\to 0^{+}$.
  • Each of the limiting profiles shown demonstrate that solutions approach equilibrium for $a$ large enough. In each case, the transition from nontrivial limiting behavior to extinction is sudden, and occurs via a saddle-node bifurcation (either a saddle-node of limit cycles or of equilibria). Thus, in a situation where a predator-prey system is near the extinction threshold $a_{\text{ext}}$, a small perturbation of the exclusion zone size can, without warning, lead to the catastrophic extinction of both species. This is a particularly alarming possibility for the scenario shown in Figure 5 because the optimal value of $a$ for the predators is located so precariously close to the extinction threshold.
@@
7. Discussion
We have explored a diffusive predator-prey model with an exclusion zone where the predators cannot go. This model is relevant to a variety of ecological scenarios, such as the development of marine protected areas (MPAs) where fishing is excluded from certain areas [8, 16], landscapes where property boundaries allow small-bodied species to pass but not larger ones [33], and situations where predator territoriality and movement behaviors concentrate their activity spatially [38, 17]. A similar model was studied by [13], but here we expand on that work to consider scenarios in which the prey species is subject to an Allee effect [11, 50]. More specifically, we considered what ecologists term a strong Allee effect in which the growth rate of the prey population is negative for a range of small densities [50]. Review papers characterize the biological evidence for Allee effects in natural and experimental settings [25, 40].
We considered versions of this diffusive predator-prey model for both one and multiple spatial dimensions. We proved extinction of both predator and prey if the exclusion zone $\Omega\setminus A$ is small and $\gamma$ (the predator’s per capita mortality rate in the absence of prey) is sufficiently small, and we have analyzed the limited conditions for coexistence when there is no predator exclusion zone (i.e., when $L=a$).
Holding other parameters fixed, the conditions for coexistence of predator and prey are somewhat relaxed when $\Omega\setminus A$ is sufficiently large, demonstrating the value of the exclusion zone. Interestingly, the transition from coexistence to mutual extinction can be quite abrupt as seen in Figures 4-6. Specifically, for scenarios in which $a$ is close to $L$ (i.e., when the exclusion zone is small), a discontinuity exists in which the long-term average population densities of both predators and prey transition from positive values to zero as $a$ passes through a threshold. This sharp transition would be a concern in the context of MPAs because it implies that fisheries could collapse quickly, without much warning, as a consequence of changes in their spatial management. That is, merely changing the location of the boundary of an MPA could be sufficient to drive a well-functioning fisher-fisheries system to collapse, if that boundary change resulted in the MPA being too small. A similarly abrupt collapse to mutual extinction would occur if the MPA were just barely big enough to facilitate coexistence, but some subset of rule-breaking fishers decided to violate the boundary and fish inside the exclusion zone (for real-world examples, see [22, 2]). In addition, an abrupt population collapse could also occur in the related scenario in which the boundary of the exclusion zone remained fixed, but some exogenous factor (e.g., a feature of global change such as water temperature) reduced the size of the habitable domain such that the MPA is no longer sufficiently large [10, 48].
@@ 7. Discussion
Overall, we have found that a diffusive predator-prey model with an exclusion zone limiting the predator and an Allee effect for the prey is capable of a rich array of dynamics. Beyond the “usual suspects” of two species coexistence and Hopf bifurcations to periodic predator-prey cycles, we also found that the model can exhibit a thin-limit with a positive predator population size when the exclusion zone is very large and sudden, catastrophic collapse from coexistence to mutual extinction if the exclusion zone becomes too small. Lastly, we found it especially interesting that the complex outcomes for this model were achieved largely through “spatial changes” traceable to alteration of the size of the predation zone and the predator-free exclusion zone rather than to the addition of ever more complex expressions for population growth, functional responses, and modes of dispersal. Some of this complexity was, of course, only possible through “spatial changes” in combination with the nonlinearity induced by the Allee effect, but the fact remains that assumptions about spatial structure can matter quite strongly, even in a model as well studied as the Lotka-Volterra predator-prey system.
7.1. Future work
@@
Third, our computations indicate that the system supports a wide variety of bifurcations, including Hopf and saddle-node bifurcations of equilibria and limit cycles, and possibly a period-doubling route to chaos. An analytical treatment these bifurcations would substantially deepen our understanding of the complexity introduced to a predator–prey system by the presence of an exclusion zone.
@@
In summary, this study establishes conditions for coexistence in predator-prey systems with exclusion zones and the Allee effect, showing how spatial structure can sustain or disrupt the survival of both species. Our analytical and numerical results together reveal rich behaviors shaped by domain geometry and diffusion, and lay the groundwork for future studies.

### v2

@@ 5.1 Computations of equilibria 1D
    This indicates that population density blows up as $a\to 0^{+}$ while total population approaches a finite limit. This supports the thin-limit analysis of Section 4. In fact, the formula (48) allows us to directly calculate the limit of $\overline{V}(a)$ and its derivative as $a\to 0^{+}$.
  • Each of the limiting profiles shown demonstrate that solutions approach equilibrium for $a$ large enough. In each case, the transition from nontrivial limiting behavior to extinction is sudden, and appears to occur via a saddle-node bifurcation (either a saddle-node of limit cycles or of equilibria). Thus, in a situation where a predator–prey system is near the extinction threshold $a_{\text{ext}}$, a small perturbation of the exclusion zone size can, without warning, lead to the catastrophic extinction of both species. This is a particularly alarming possibility for the scenario shown in Figure 5 because the optimal value of $a$ for the predators is located so precariously close to the extinction threshold.
@@
6 Discussion
We have explored a diffusive predator–prey model with an exclusion zone where the predators cannot enter. This model is relevant to a variety of ecological scenarios, such as the development of marine protected areas (MPAs) where fishing is excluded from certain areas [7, 14], landscapes where property boundaries allow small-bodied species to pass but not larger ones [32], and situations where predator territoriality and movement behaviors concentrate their activity spatially [37, 15]. A similar model was studied by [13], but here we expand on that work to consider scenarios in which the prey species is subject to an Allee effect [10, 49]. More specifically, we considered what ecologists term a strong Allee effect in which the growth rate of the prey population is negative for a range of small densities [49]. Review papers characterize the biological evidence for Allee effects in natural and experimental settings [24, 39].
Holding other parameters fixed, the conditions for coexistence of predator and prey are somewhat relaxed when $\Omega\setminus A$ is sufficiently large, demonstrating the value of the exclusion zone. Interestingly, the transition from coexistence to mutual extinction can be quite abrupt as seen in Figures 4-6. Specifically, for many scenarios in which $a$ is close to $L$ (i.e., when the exclusion zone is small), we observe an apparent discontinuity in which the long-term average population densities of both predators and prey transition from positive values to zero as $a$ passes through a threshold. This sharp transition would be a concern in the context of MPAs because it implies that fisheries could collapse quickly, without much warning, as a consequence of changes in their spatial management. That is, merely changing the location of the boundary of an MPA could be sufficient to drive a well-functioning fisher-fisheries system to collapse, if that boundary change resulted in the MPA being too small. A similarly abrupt collapse to mutual extinction would occur if the MPA were just barely big enough to facilitate coexistence, but some subset of rule-breaking fishers decided to violate the boundary and fish inside the exclusion zone (for real-world examples, see [21, 2]). In addition, an abrupt population collapse could also occur in the related scenario in which the boundary of the exclusion zone remained fixed, but some exogenous factor (e.g., a feature of global change such as water temperature) reduced the size of the habitable domain such that the MPA is no longer sufficiently large [9, 47].
@@ 6 Discussion
Overall, we have found that a diffusive predator–prey model with an exclusion zone limiting the predator and an Allee effect for the prey is capable of a rich array of dynamics. Beyond the “usual suspects” of two species coexistence and Hopf bifurcations to periodic predator–prey cycles, we also found that the model can exhibit a thin-limit with a positive predator population size when the exclusion zone is very large and sudden, catastrophic collapse from coexistence to mutual extinction if the exclusion zone becomes too small. Lastly, we found it especially interesting that the complex outcomes for this model were achieved largely through “spatial changes” traceable to alteration of the size of the predation zone and the predator-free exclusion zone rather than to the addition of ever more complex expressions for population growth, functional responses, and modes of dispersal. Some of this complexity was, of course, only possible through “spatial changes” in combination with the nonlinearity induced by the Allee effect, but the fact remains that assumptions about spatial structure can matter quite strongly, even in a model as well studied as the Lotka-Volterra predator–prey system.
6.1 Future work
@@
Third, our computations indicate that the system may support a wide variety of bifurcations, including Hopf and saddle-node bifurcations of equilibria and limit cycles, and possibly a period-doubling route to chaos. An analytical treatment these bifurcations would substantially deepen our understanding of the complexity introduced to a predator–prey system by the presence of an exclusion zone.
@@
In summary, this study establishes conditions for coexistence in predator–prey systems with exclusion zones and the Allee effect, showing how spatial structure can sustain or disrupt the survival of both species. Our analytical and numerical results together reveal rich behaviors shaped by domain geometry and diffusion, and lay the groundwork for future studies.
Declarations
Funding. W. F. Fagan was supported by NSF DMS-2451241 The Mathematics of Animal Migration.
Conflict of Interest. The authors declare that they have no conflicts of interest or competing interests associated with this work.
Data Availability. No external data were used in this manuscript. Data generated in this manuscript (in particular, the data shown in the figures in Section 5) are available upon request. Additionally, the software used in Section 5 to solve the one dimensional model is available at https://github.com/alexsafsten/ExclusionZone1D.
Appendix A Proof of Proposition 4.3
Proof.
By Proposition 4.1 we may select $b_{1}$ so that if $L>b_{1}$, then $\zeta_{0,L}(L)>\theta^{\prime}$. We will show that there exists $b_{0}\geq b_{1}$ such that (30) has a unique solution $u$ whenever $L>L_{0}$ such that $\zeta_{0,L}<u<1$. Then letting $\eta_{0}=1-\zeta_{0,b_{0}}(b_{0})$, the existence and uniqueness aspect of the result is proved. The fact that this solution is also monotone and $u(L)$ is increasing with $u(L)\to 1$ as $L\to\infty$ will follow from the proof of uniqueness.
Observe that $\zeta_{0,L}$ is a subsolution of (30) and the constant 11 is a supersolution. Therefore, there exists a solution $u$ with $\zeta_{0,L}<u<1$. It remains to show that, if $L$ is sufficiently large, this is the only solution such that $\zeta_{a_{0},L}<u<1$. We begin by showing that such a solution is monotone increasing. Multiplying (30) by $u^{\prime}$ and integrating, we find that $u$ satisfies the conservation law (28), so for any $x$,
$\frac{d_{u}}{2}(u^{\prime}(x))^{2}+F(u(x))=F(u(L)).$ (64)
Since $u(L)>\zeta(L)>\theta^{\prime}$, $F(u(L))>0$. Whenever $F$ is positive, it is one-to-one, so we conclude that if $u^{\prime}(x)=0$ for any $x$, then $u(x)=u(L)$. Therefore, $u$ cannot achieve an internal maximum at any point $x_{0}\in(0,L)$ because at this maximum, we would have $u(x_{0})=u(L)$, but then either $u$ is constant on $x_{0},L]$ or $u$ achieves a minimum at a point $x_{1}\in(x_{0},L)$ where $u(x_{1})<u(L)$. Neither of these is possible. Similarly, $u$ cannot achieve a local minimum. We conclude that $u$ is monotone. Since $u(L)>\theta$, $u^{\prime}(L)=0$, and $d_{u}u^{\prime\prime}=f(u)$, we have $u^{\prime\prime}(x)<0$ for $x$ near $L$ and thus, $u^{\prime}(x)>0$ for such $x$. But since $u$ is monotone, we must therefore have $u^{\prime}(x)>0$ for all $x\in[0,L)$.
Since $u$ is increasing, we can use (64) to write $u^{\prime}$ as:
$u^{\prime}=\sqrt{\frac{2}{d_{u}}\left(F(u(L))-F(u)\right)}$ (65)
Again, due to the strict monotonicity of $u$, we may invert this differential equation to compute $x$ as a function of $u$. Then, letting $q=u(L)$ and $p=\gamma/\alpha$, we may write $L$ as a function of $q$:
$L=\mathcal{L}(q)=\sqrt{\frac{d_{u}}{2}}\int_{p}^{q}\frac{1}{\sqrt{F(q)-F(u)}}\,du.$ (66)
Thus, (30) has a monotone increasing solution $u$ on $[0,L]$ such that $u(L)=q<1$ precisely when $L=\mathcal{L}(q)$. If there were two such solutions $u_{1}$ and $u_{2}$, we would have $u_{1}(L)=q_{1}$ and $u_{2}(L)=q_{2}$. If $q_{1}=q_{2}$, then by uniqueness of solutions to initial value problems, $u_{1}=u_{2}$. Otherwise, we have $\mathcal{L}(q_{1})=\mathcal{L}(q_{2})$. We will show that for $q$ sufficiently close to 11, $\mathcal{L}(q)$ is strictly increasing, establishing uniqueness.
Let $\nu=-f^{\prime}(1)>0$ and let $\delta>0$ be fixed such that both of the following hold for $1-2\delta<s<1$:
  • $2\nu\leq f^{\prime}(s)\leq-\nu/2$
  • $\frac{\nu}{2}(1-s)<f(s)<2\nu(1-s)$
We also assume $1-2\delta>\theta^{\prime}$. Suppose $q>1-\delta$. Then we break up (66) into two integrals:
$\mathcal{L}(q)$ $=A(q)+B(q)$
$=\sqrt{\frac{d_{u}}{2}}\int_{p}^{q-\delta}\frac{1}{\sqrt{F(q)-F(u)}}\,du+\sqrt{\frac{d_{u}}{2}}\int_{q-\delta}^{q}\frac{1}{\sqrt{F(q)-F(u)}}\,du.$
Since $q>\theta^{\prime}$, there is no $u\leq q-\delta$ such that $F(u)=F(q)$. Then the function $(u,q)\mapsto F(q)-F(u)$ is continuous and positive on the compact set
$T=\{(u,q)\in\mathbb{R}^{2}:1-\delta\leq q\leq 1,\;p\leq u\leq q-\delta\},$
so we conclude that there exists $m>0$ so that $F(q)-F(u)\geq m$ for all $1-\delta\leq q\leq 1$ and $p\leq u\leq q-\delta$. Thus,
$|A(q)|\leq\sqrt{\frac{d_{u}}{2}}\frac{(1-p)}{\sqrt{m}}.$
and
$|A^{\prime}(q)|$ $=\sqrt{\frac{d_{u}}{2}}\left|\frac{1}{\sqrt{F(q)-F(q-\delta)}}-\frac{F^{\prime}(q)}{2}\int_{p}^{q-\delta}\frac{1}{(F(q)-F(u))^{3/2}}\,du\right|$
$\leq\sqrt{\frac{d_{u}}{2}}\left(\frac{1}{\sqrt{m}}+\frac{\|f\|_{C^{0}}}{2}\frac{(1-p)}{m^{3/2}}\right).$
Therefore, both $A(q)$ and $A^{\prime}(q)$ are uniformly bounded as $q\to 1^{-}$.
Now we show $\lim_{q\to 1^{-}}B(q)=\lim_{q\to 1^{-}}B^{\prime}(q)=\infty$ so for $q$ sufficiently close to 11, $\mathcal{L}(q)$ is increasing and unbounded. We may write
$B(q)=\sqrt{\frac{d_{u}}{2}}\int_{0}^{\delta}\frac{1}{\sqrt{F(q)-F(q-s)}}\,ds.$
That $B(q)$ is differentiable is nontrivial because the integrand is unbounded. Therefore, we first show that $B(q)$ is differentiable and that the expression for the derivative can be found by differentiating under the integral sign, as expected.
If $B^{\prime}(q)$ exists, it is the limit of the following finite difference as $h\to 0$:
$\frac{B(q+h)-B(q)}{h}=\sqrt{\frac{d_{u}}{2}}\int_{0}^{\delta}\frac{1}{h}\left(\frac{1}{\sqrt{F(q+h)-F(q+h-s)}}-\frac{1}{\sqrt{F(q)-F(q-s)}}\right)\,ds.$ (67)
We will show that there exists a constant $C$ such that for $|h|$ sufficiently small, the integrand of (67) is bounded by $C/\sqrt{s}$, which is integrable. Therefore, by the Lebesgue Dominated Convergence theorem, we may pass the limit into the integral.
First, we note that
$F(q)-F(q-s)=\int_{q-s}^{q}f(s^{\prime})\,ds^{\prime},$ (68)
so we may bound this quantity for $1-2\delta<q-s<q<1$:
$\frac{\nu}{4}s(s+2\varepsilon)=\frac{\nu}{2}\int_{q-s}^{q}(1-s^{\prime})\,ds^{\prime}\leq F(q)-F(q-s)\leq 2\nu\int_{q-s}^{q}(1-s^{\prime})\,ds^{\prime}=\nu s(s+2\varepsilon),$ (69)
where $\varepsilon=1-q$. Thus, the integrand in (67) is bounded as follows, assuming $|h|<\varepsilon/2$:
$\left|\frac{1}{h}\left(\frac{1}{\sqrt{F(q+h)-F(q+h-s)}}-\frac{1}{\sqrt{F(q)-F(q-s)}}\right)\right|$
$=\frac{|\sqrt{F(q)-F(q-s)}-\sqrt{F(q+h)-F(q+h-s)}|}{h\sqrt{F(q)-F(q-s)}\sqrt{F(q+h)-F(q+h-s)|}}$
$=\frac{8|F(q)-F(q-s)-F(q+h)-F(q+h-s)|}{h\left((F(q)-F(q-s))\sqrt{F(q+h)-F(q+h-s)}+(F(q+h)-F(q+h-s))\sqrt{F(q)-F(q-s)}\right)}$
$\leq\frac{8\left|\int_{q-s}^{q}\int_{y+h}^{y}f^{\prime}(z)\,dz\,dy\right|}{h(\nu s)^{3/2}\left((s+2\varepsilon)\sqrt{s+2(\varepsilon-h)}+(s+2(\varepsilon-h))\sqrt{s+2\varepsilon}\right)}$
$\leq\frac{16\nu hs}{h(\nu s\varepsilon)^{3/2}\left(2+\sqrt{2}\right)}$
$\leq\frac{C}{\sqrt{s}},\quad\text{where}\quad C=\frac{16}{\sqrt{\nu\varepsilon^{3}}(2+\sqrt{2})}.$
Thus, $B(q)$ is differentiable and
$B^{\prime}(q)=\frac{1}{2}\sqrt{\frac{d_{u}}{2}}\int_{0}^{\delta}\frac{F^{\prime}(q-s)-F^{\prime}(q)}{(F(q)-F(q-s))^{3/2}}\,ds.$
We may formulate a similar bound to (69) for $F^{\prime}$:
$F^{\prime}(q-s)-F^{\prime}(q)=-\int_{q-s}^{q}f^{\prime}(r)\,dr\geq\frac{\nu s}{2}.$
Therefore, we calculate
$\begin{split}B(q)&\geq\sqrt{\frac{d_{u}}{2\nu}}\int_{0}^{\delta}\frac{1}{\sqrt{s(s+2\varepsilon)}}\,ds=\left.-\sqrt{\frac{2d_{u}}{\nu}}\log\left(\sqrt{s+2\varepsilon}-\sqrt{s}\right)\right|_{0}^{\delta}=\sqrt{\frac{2d_{u}}{\nu}}\operatorname{arcsinh}\left(\sqrt{\frac{\delta}{2\varepsilon}}\right),\end{split}$ (70)
and furthermore,
$B^{\prime}(q)\geq\frac{1}{4}\sqrt{\frac{d_{u}}{2\nu}}\int_{0}^{\delta}\frac{1}{\sqrt{s}(s+2\varepsilon)^{3/2}}\,ds=\left.\frac{1}{4}\sqrt{\frac{d_{u}}{2\nu}}\frac{\sqrt{s}}{\varepsilon\sqrt{s+2\varepsilon}}\right|_{0}^{\delta}=\frac{1}{4\varepsilon}\sqrt{\frac{d_{u}}{2\nu}}\sqrt{\frac{\delta}{\delta+2\varepsilon}}.$ (71)
The expressions (70) and (71) show that $B(q)$ and $B^{\prime}(q)$ both approach $\infty$ as $q\to 1^{-}$ (which corresponds to $\varepsilon\to 0^{+}$). Since $A(q)$ and $A^{\prime}(q)$ are both bounded, we conclude that there exists $q_{0}$ with $1-\delta<q_{0}<1$ so that $L(q)$ is increasing for $q_{0}<q<1$ and moreover, $L(q)>L(q^{\prime})$ for each $q^{\prime}<q_{0}$. Let $b_{0}=\max\{\mathcal{L}(q_{0}),b_{1}\}$ and $\eta_{0}=\zeta_{0,b_{0}}$. Then for all $L>b_{0}$, (30) has a unique solution $u$ so that $1-\eta_{0}<u(L)<1$. Moreover, we have already shown that this solution is monotone increasing, and that as $L\to\infty$, $u(L)$ must monotonically approach 11. ∎